If a 25.0 L sample of gas at C is heated, what is the new sample of the gas at constant pressure?

1
Expert's answer
2021-05-21T02:33:45-0400

According to gas law

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"V_2=\frac{V_1×T_2}{T_1}=\frac{25×298}{273}=27.3l"